1. During a given year, the following activities occur in a simple economy consisting of a miller and a baker The miller pays his workers $ 250,000 to mill 120 tons of flour. The flour is then sold to a baker for $ 530,000. The baker pays his workers 340,000 to make 750,000 loaves of bread, which she then sells directly to consumers for $ 1,000,000. a. Using the "final goods" approach, what is the GDP in this economy? b. What is the value added at each stage of production? Using the "value added" approach, what is GDP? c. What are the total wages and profits earned? Using the income approach, what is GDP? 2. Consider a simple economy that produces only 3 products. Use the in- formation in the following table to calculate the growth rate of real GDP from 2018 to 2019. The base year is 2010. 2018 2019 2010 Price Price Product Price Q $ 10.00 3$8.00 $11.00 Haircuts 7.45 7.99 Hamburgers 10 7.00 16 17 1225.00 18 25.00 20 22.0 Books 3. You are an employee of the Bureau of Labor Statistics involved in the monthly survey of households used to estimate the unemployment rate In each of the following cases, classify the individual as ployed employed, or not in the labor force. Explain you classification. unem-
a. During the entire week containing the 12th of the month, Joseph misses work simply because he didn't feel like going in to work. b. Rachel is a 22-year-old college student, who is out of school for the summer, is looking for a job. c. Carley is a hugely successful film star, age 12, who has earned over $10,000,000 each year for the past five years d. Andy is a 32-year-old man with a Ph.D. in economics. He has not been able to find a teaching position and is driving a cab 30 hours a week e. Susan is a 70-year-old woman who left her job to help her son in his business for 10 hours a week. 4. Using the table below, calculate the unemployment rate Structural Unemployment Frictional Unemployment Cyclical Unemployment Employed Discouraged Workers 400 200 100 8,000 500 5. An economy produces three goods: cars, computers, and oranges. Pro- duction units and prices per unit for years 2018 and 2019 are as follows: 2019 2018 Quantity Price($) | Quantity Price($) Cars 12 25,000 10 20,000 1,000 60 750 40 Computers Oranges 1 1,000 1 1,000 Answer the following: a. Using 2018 as the base year, what is the GDP deflator for 2018 and 2019? What is the rate of inflation over this period?
b. Repeat a. using 2019 as the base year c. Does the choice of base year affect the rate of inflation computed using the GDP deflator? d. Calculate the rate of inflation for 2018 and 2019 using the CPI and 2018 as the base year
Answer #1

Answer 1)

a) The final goods approach is calculated by adding all final expenditures.

GDP = Consumption + Investment expenditure + Government expenditure + Net Exports

GDP = 1,000,000 + 0 + 0 + 0

GDP = 1,000,000

Since , there is no investment and government expenditure so it is equal to zero. As well as there is no import and export of product across countries so net exports will be zero. However there is consumption of $1,000,000, when baker sells directly to the consumers.

b) In the value added approach , only the final value of good is included in total output , to avoid an issue of double counting. When the miller sells product to baker the value added is $530,000. But this is not the final value of good produced as it is an intermediate good for baker from which he makes bread. After the final good is produced , it is sold to the consumers. Thus , the final value of good , bread , will be $1,000,000. In other words, GDP = $1,000,000 in value added approach.

c) Wages paid by miller = $ 250,000 and Wages paid by baker = $340,000

Thus , total wages paid = 250,000 + 340,000 = $590,000

Profits earned by miller = flour sold to baker - wages paid by miller

Profits earned by miller = 530,000 - 250,000 = $ 280,000

Profits earned by baker = Bread sold to consumers - ( flour purchased by baker + wages paid by baker )

profits earned by miller =1,000,000 - (530,000 + 340,000 ) = 1,000,000 - 870,000 = $130,000

Total profits earned = profits of miller + profits of baker = 280,000 + 130,000 = $410,000

Therefore, according to the question ,In income approach

GDP = Wages + Profits = 590,000 + 410,000 = $1,000,000